# FlexDiT: Dynamic Token Density Control for Diffusion Transformer

### v1

@@ Abstract
Diffusion Transformers (DiT) deliver impressive generative performance but face prohibitive computational demands due to both the quadratic complexity of token-based self-attention and the need for extensive sampling steps. While recent research has focused on accelerating sampling, the structural inefficiencies of DiT remain underexplored. We propose FlexDiT, a framework that dynamically adapts token density across both spatial and temporal dimensions to achieve computational efficiency without compromising generation quality. Spatially, FlexDiT employs a three-segment architecture that allocates token density based on feature requirements at each layer: Poolingformer in the bottom layers for efficient global feature extraction, Sparse-Dense Token Modules (SDTM) in the middle layers to balance global context with local detail, and dense tokens in the top layers to refine high-frequency details. Temporally, FlexDiT dynamically modulates token density across denoising stages, progressively increasing token count as finer details emerge in later timesteps. This synergy between FlexDiT’s spatially adaptive architecture and its temporal pruning strategy enables a unified framework that balances efficiency and fidelity throughout the generation process. Our experiments demonstrate FlexDiT’s effectiveness, achieving a 55% reduction in FLOPs and a 175% improvement in inference speed on DiT-XL with only a 0.09 increase in FID score on 512$\times$512 ImageNet images, a 56% reduction in FLOPs across video generation datasets including FaceForensics, SkyTimelapse, UCF101, and Taichi-HD, and a 69% improvement in inference speed on PixArt-$\alpha$ on text-to-image generation task with a 0.24 FID score decrease. FlexDiT provides a scalable solution for high-quality diffusion-based generation compatible with further sampling optimization techniques. Code will be available at https://github.com/changsn/FlexDiT.
@@ 1 Introduction
Diffusion models [19, 12, 43, 3] have emerged as powerful tools in visual generation, producing photorealistic images and videos by gradually refining structured content from noise. Leveraging the scalability of Transformers, Diffusion Transformer (DiT) models [40, 2] extend these capabilities to more complex and detailed tasks [13, 62] and form the backbone of advanced generation frameworks such as Sora [6]. Yet, despite their impressive generative capabilities, DiT models face significant computational limitations that restrict their broader applicability, especially in scenarios where efficiency is paramount.
@@
The computational burden of DiT models arises mainly from the need for numerous sampling steps in the denoising process and the quadratic complexity of Transformer structures with respect to the number of tokens. Although recent research has focused on reducing this burden by accelerating the sampling process through techniques such as ODE solvers [51, 29], flow-based methods [26, 28], and knowledge distillation [53, 46, 31, 34, 63], these approaches often overlook the core architectural inefficiencies of DiT itself. Unlike U-Net-based architectures [44], which mitigate complexity through a contracting-expansive structure, DiT’s reliance on token-level self-attention incurs high costs that scale with model size and token count. This bottleneck highlights a critical need for DiT-specific innovations that manage token density intelligently, balancing efficiency and quality without sacrificing generation fidelity.
@@
To address this challenge, we examine the distribution of attention across DiT’s layers and sampling steps, seeking to uncover structural patterns that could guide more efficient token management. As illustrated in Figure 1(a), attention maps reveal that tokens capture different levels of granularity across layers: in the bottom layers, such as layer 0, attention maps appear uniformly distributed, indicating a focus on broad, global features akin to global pooling. Meanwhile, middle layers alternate in their attention, with certain layers (e.g., layer 6 and layer 26) capturing local details, while others (e.g., layer 22) emphasize global structure. This observation is quantified in Figure 1(b), where the variance in attention scores highlights DiT’s consistent pattern of alternation between global and local feature extraction across layers. In addition to spatial analysis, Figure 1(b) also reveals insights across sampling steps. We observe that, while the pattern of alternation between global and local focus remains stable, attention variance increases as the denoising process advances, indicating a growing emphasis on local information at lower-noise stages. This analysis yields three core insights: (1) the initial self-attention layers exhibit minimal variance, showing low discrepancy in feature extraction across tokens; (2) DiT’s architecture inherently alternates between global and local feature focus across layers, a pattern consistent across all sampling steps; and (3) as denoising progresses, the model increasingly prioritizes local details, dynamically adapting to heightened demands for detail at later stages.
@@
Building on these insights, we propose FlexDiT, which reimagines token density management as a dynamic process adapting across both spatial layers and temporal stages. Spatially, shallow layers capture smooth, global features, making complex self-attention less efficient. Subsequent layers alternate between high-frequency local details and low-frequency global information, where sparse tokens efficiently capture global features, and dense tokens refine local details. Temporally, as denoising advances, the need for local detail increases, motivating a dynamic token adjustment strategy. This dual-layered adaptation integrates layer-wise token modulation—balancing efficiency and detail within each sampling step—and a timestep-wise pruning strategy that adjusts token density dynamically. Together, these spatial and temporal adaptations allow FlexDiT to achieve computational efficiency without sacrificing high-fidelity detail, as reflected in FlexDiT’s architectural design and pruning strategy, tailored to meet the unique spatial and temporal demands of the generation process.
@@
FlexDiT Architecture: FlexDiT’s architecture employs a three-segment design that aligns token density with the features each layer captures. In the bottom layers, we replace self-attention with a poolingformer structure to capture broad global features through average pooling, reducing computation while preserving essential structure. In the middle layers, FlexDiT introduces Sparse-Dense Token Modules (SDTM), blending sparse tokens for global structure with dense tokens to refine local details. This design enables FlexDiT to retain efficiency while maintaining stability and detail. In the top layers, the model processes all tokens densely, focusing on high-frequency details for refined output quality.
@@
Timestep-Wise Pruning Rate Strategy: Complementing spatial adjustments, FlexDiT’s timestep-wise pruning rate strategy adapts token density across denoising stages. In early stages, where broad, low-frequency structures dominate, FlexDiT applies a high pruning rate, conserving resources. As the process progresses and intricate details emerge, the pruning rate decreases, allowing for a gradual increase in token density. This adjustment aligns computational effort with each stage’s complexity, dynamically balancing efficiency and detail to optimize resources for high-fidelity output.
@@
Our empirical results substantiate the effectiveness of FlexDiT’s integrated approach. FlexDiT reduces the FLOPs of DiT-XL by 55% and improves inference speed by 175% on 512$\times$512 ImageNet [11] images, with only a increase of 0.09 in FID score [17]. On the Latte-XL dataset [33], FlexDiT achieves a 56% reduction in FLOPs across standard video generation datasets, including FaceForensics [45], SkyTimelapse [60], UCF101 [54], and Taichi-HD [50]. Additionally, on the more challenging text-to-image generation task, we achieve a 69% improvement in inference speed on PixArt-$\alpha$ with a 0.24 FID score reduction. These results demonstrate FlexDiT’s capability to provide an efficient, high-quality architecture for diffusion-based generation, compatible with further sampling optimization techniques for enhanced efficiency.
(a) Architecture of FlexDiT.
@@
Figure 2: Architecture of FlexDiT and timestep-wise pruning rate strategy. The architecture consists of three segments: bottom, middle, and top. The bottom segment includes poolingformers (PF). The middle segment comprises multiple sparse-dense token modules (SDTMs), where sparse token generation transformers (“Sparse Gen”) and dense token recovery transformers (“Dense Rec”) alternate to balance global and detailed information processing. The top segment contains standard transformers (“Tr”) for final processing. During the denoising process, we apply varying pruning rates $r$ at different stages (depicted as four stages).
@@ 2 Related works
Although diffusion models [19, 43] have achieved impressive performance in generation tasks, their generation speed remains a significant bottleneck for broader applications. This limitation arises primarily from the step-wise reverse diffusion process and the high computational cost associated with single model evaluations.
@@
This direction focuses on optimizing sampling steps in diffusion models. Some approaches [51, 29, 1, 30] design new solvers to achieve faster sampling with fewer steps. Consistency models (CMs) [53, 52, 31, 58, 21] are closely related to diffusion models, achieving by distilling pre-trained diffusion models or training from scratch. These models learn a one-step mapping between noise and data, and all the points of the sampling trajectory map to the same initial data point. Flow-based [26, 28, 67, 13, 68] approaches straighten the transport trajectories among different distributions and thereby realize efficient inference. Knowledge distillation are applied by some methods [53, 46, 31, 34, 63] to reduce the sampling steps and match the output of the combined conditional and unconditional models. Parallel sampling techniques [49, 66, 56] employ Picard-Lindelöf iteration or Fourier neural operators for solving SDE/ODE.
@@
Another avenue focuses on improving inference time within a single model evaluation. Some methods attempt to compress the size of diffusion models via pruning [14] or quantization [25, 48]. Spectral Diffusion [61] boosts structure design by incorporating frequency dynamics and priors. OMS-DPM [27] proposes a model schedule that selects small models or large models at specific sampling steps to balance generation quality and inference speed. Deepcache [32] reuse or share similar features across sampling steps. The early stopping mechanism in diffusion is explored in  [24, 36, 55].
Most of the above methods are designed for general diffusion models. Recently, the Diffusion Transformer (DiT) has emerged as a more potential backbone, surpassing the previously dominant U-Net architectures. Unlike U-Net-based models, DiT’s inefficiency stems largely from the self-attention mechanism. Limited research has specifically addressed the efficiency of DiT-based models.  [42] identifies the redundancies in the self-attention and adopts timestep dynamic mediator tokens to compress attention complexity, achieving some performances but limited FLOPs reduction, as this approach primarily reduces the number of keys and values rather than the overall tokens count. In Vision transformers for classification task, numerous methods [59, 9, 15, 7, 5, 69, 35] successfully removed redundant tokens to improve performance-efficiency trade-offs. ToMeSD [4] is the first to explore token reduction in diffusion models. However, its performance on DiT was suboptimal, as shown in  [36, 65]. U-Net can be regarded as a sparse network due to its contracting and expansive structure. EDT [10] applies a novel architecture similar to U-Net, while they still have 1.4 FID score gap compared to its baseline model MDTv2 [16] via 2, [^0] K iterations. These findings indicate that directly transferring token reduction techniques from classification tasks or reusing U-Net structures may be insufficient for DiT. Instead, a specialized token reduction strategy is needed.
We analyze the structure of DiT and observe that local and global features are captured alternately across layers. Inspired by this insight, FlexDiT introduces a novel token reduction strategy that decouples the generation process into separate integration of global and local information, leveraging sparse tokens for global feature extraction and dense tokens for local detail refinement. By alternating between sparse and dense tokens, FlexDiT ensures high efficiency and performance.
@@ 3 Method
We first provide an overview of diffusion models and DiT in Section 1. FlexDiT architecture is then introduced in Section 3.2, followed by our timestep-wise pruning rate strategy in Section 6. Finally, Section 3.4 details the initialization and fine-tuning of our network.
3.1 Preliminary
Diffusion models
generate visual contents (such as images and vidoes) from random noise through a series of denoising steps. These models typically consist of a forward diffusion process and a reverse denoising process. In the forward process, given an visual input $x_{0}\sim q(x)$ sampled from the data distribution, Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ is progressively added over $T$ steps. This process is defined as $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$, where $t$ and $\beta_{t}$ denote the sampling step and noise schedule, separately. In the reverse process, the model removes the noise and reconstructs $x_{0}$ from $x_{T}\sim\mathcal{N}(0,I)$ using $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sum_{\theta}(x_{t},t))$, where $\mu_{\theta}(x_{t},t)$ and $\sum_{\theta}(x_{t},t)$ represent the mean and variance of the Gaussian distribution.
Diffusion transformer (DiT)
exhibits the scalability and promising performance of transformers. Similar to ViT, DiT consists of layers composed of a multi-head self-attention (MHSA) layer and a multi-layer perceptron (MLP) layer, described as
$\begin{gathered}X=X+\alpha MHSA(\gamma X+\beta),\\ X=X+\alpha^{\prime}MLP(\gamma^{\prime}X+\beta^{\prime}),\end{gathered}$ (1)
where $X\in\mathbb{R}^{N\times C}$ denotes input tokens, $N$ is the number of tokens, and $C$ is the channel dimension. The parameters $\{\alpha,\gamma,\beta,\alpha^{\prime},\gamma^{\prime},\beta^{\prime}\}$ are produced by an adaptive layer norm (AdaLN) block [41], which takes the condition embedding and sampling step embedding as inputs.
3.2 FlexDiT
@@
The FlexDiT architecture is illustrated in Figure 2(a), where the transformer layers of DiT based model are divided into three groups: bottom, middle, and top. The bottom layers leverage poolingformers to efficiently capture global features. The middle layers contain multiple sparse-dense token modules (SDTM), which decouple the representation process into global structure capture and local details enhancement using sparse and dense tokens, respectively. The top layers retain the standard transformers, processing dense tokens to generate the final predictions at each sampling step. The primary computational savings are achieved through sparse tokens, hence the majority of transformer layers are located in the middle section to maximize efficiency.
@@
(b) Replacing the first two attention maps in DiT-XL with full-one matrices without fine-tuning.
Figure 3: Comparison of images generated by the standard DiT-XL model and a modified DiT-XL model where the first two attention maps are replaced with full-one matrices. Both sets of images exhibit similar visual quality.
@@
As shown in the first subfigure in Figure 1(a), attention scores in the bottom layers are uniformly distributed, with each token evenly extracting global features, akin to a global average pooling operator. To investigate further, we conducted a toy experiment where the attention maps in the first two transformer layers were replaced with matrices filled with $1$, without any fine-tuning. Figure 3 shows that, using the same initial noise and random seed, the images generated by both the original DiT-XL and the modified DiT-XL are nearly identical, suggesting that complex self-attention calculations offer limited additional information. Given that self-attention in the bottom layers captures global features, we question whether sparse tokens can be used in these layers. Extensive experiments, however, demonstrated that applying sparse tokens in the lower layers leads to unstable training; maintaining a full set of tokens in these layers is essential.
Based on above analysis, we adopt poolingformers to replace the original transformers. In poolingformer, we remove the queries and keys as attention maps are not computed. Instead, we perform global pooling on the value $V\in\mathbb{R}^{N\times C}$ and integrate it into input tokens $X$, which can be represented as
@@
where $\bar{V}$ is the mean along the dimension of $N$, i.e., $\bar{v}=\frac{1}{N}\sum V(\bar{v}\in\mathbb{R}^{1\times C})$, and then repeating $N$ times to shape $N\times C$. The parameters from adaLN block are omitted in all the equations for brevity. Our poolingformer can be viewed as a special case of the model in [64], behaving identically when the average pooling kernel size equals the input size.
@@
We present the sparse-dense token module (SDTM) to generate sparse and dense tokens processed by corresponding transformer layers. The high-level idea is to decouple global structure extraction and local detail extraction. Sparse tokens capture global structure information and reduce computational cost, while dense tokens enhance local detail and stabilize training. Sparse and dense tokens are converted each other within SDTM.
Initially, SDTM introduces a set of sparse tokens $X_{s}\in\mathbb{R}^{M\times C}$, where M is the number of sparse tokens, typically $M\ll N$. We define pruning rate $r=1-M/N$ to represent the sparsity degree. Sparse token interact with full-size dense tokens via an attention layer to integrate the global information:
@@
where the triplet input of $MHA$ are queries, keys, and values in turn. Sparse tokens can be initialized by adaptively pooling the dense tokens $X$. To store the structure information, we first reshape the dense tokens $X\in\mathbb{R}^{N\times C}$ into the latent shape. For instance, if the input is an image, we reshape it into shape $H\times W\times C$, then pool it across the spatial dimensions to shape $H^{\prime}\times W^{\prime}\times C$, where $M=H^{\prime}\times W^{\prime}$ and $r=1-(H^{\prime}\cdot W^{\prime})/(H\cdot W)$. The intentions of spatial pooling initialization are two folds. First, the initial sparse tokens can distribute uniformly in space and the representation of each sparse tokens is associated with a specific spatial location, beneficial for downstream tasks like image editing [20]. Second, it can prevent the semantic tokens from collapsing to one point in the following layers.
The generated sparse tokens are fed into subsequent transformer layers, referred to as sparse transformers. Since $M\ll N$, our SDTM can substantially reduce the computational cost.
At the rear of SDTM, we restore dense tokens from sparse tokens. First, we reshape the sparse tokens $X_{s}\in\mathbb{R}^{M\times C}$ into structure shape $H^{\prime}\times W^{\prime}\times C$ and upsample it to the same shape as input dense tokens $X\in\mathbb{R}^{H\times W\times C}$. We introduce two linear layers to combine upsampling sparse tokens with dense tokens, which is represented as:
@@
At the end of SDTM, several transformer layers, referred as to dense transformers process full-size dense tokens to enhance local details.
@@
Table 1: Comparison of performance between FlexDiT and DiT for class-conditional image generation task on ImageNet.
3.3 Timestep-wise pruning rate.

### v2

# SparseDiT: Token Sparsification for Efficient Diffusion Transformer

@@ Abstract
Diffusion Transformers (DiT) are renowned for their impressive generative performance; however, they are significantly constrained by considerable computational costs due to the quadratic complexity in self-attention and the extensive sampling steps required. While advancements have been made in expediting the sampling process, the underlying architectural inefficiencies within DiT remain underexplored. We introduce SparseDiT, a novel framework that implements token sparsification across spatial and temporal dimensions to enhance computational efficiency while preserving generative quality. Spatially, SparseDiT employs a tri-segment architecture that allocates token density based on feature requirements at each layer: Poolingformer in the bottom layers for efficient global feature extraction, Sparse-Dense Token Modules (SDTM) in the middle layers to balance global context with local detail, and dense tokens in the top layers to refine high-frequency details. Temporally, SparseDiT dynamically modulates token density across denoising stages, progressively increasing token count as finer details emerge in later timesteps. This synergy between SparseDiT’s spatially adaptive architecture and its temporal pruning strategy enables a unified framework that balances efficiency and fidelity throughout the generation process. Our experiments demonstrate SparseDiT’s effectiveness, achieving a 55% reduction in FLOPs and a 175% improvement in inference speed on DiT-XL with similar FID score on $\times512$ ImageNet, a 56% reduction in FLOPs across video generation datasets, and a 69% improvement in inference speed on PixArt-$\alpha$ on text-to-image generation task with a 0.24 FID score decrease. SparseDiT provides a scalable solution for high-quality diffusion-based generation compatible with sampling optimization techniques.
@@ 1 Introduction
Diffusion models ho2020denoising ; dhariwal2021diffusion ; rombach2022high ; blattmann2023stable have emerged as powerful tools in visual generation, producing photorealistic images and videos by gradually refining structured content from noise. Leveraging the scalability of Transformers, Diffusion Transformer (DiT) models peebles2023scalable ; bao2023all extend these capabilities to more complex and detailed tasks esser2024scaling ; yang2024cogvideox and form the backbone of advanced generation frameworks such as Sora brooks2024video . Yet, despite their impressive generative capabilities, DiT models face significant computational limitations that restrict their broader applicability, especially in scenarios where efficiency is paramount.
@@
The computational burden of DiT models arises mainly from the need for numerous sampling steps in the denoising process and the quadratic complexity of Transformer structures with respect to the number of tokens. Although recent research has focused on reducing this burden by accelerating the sampling process through techniques such as ODE solvers song2020denoising ; lu2022dpm , flow-based methods lipman2022flow ; liu2022flow , and knowledge distillation song2023consistency ; salimans2022progressive ; luo2023latent ; meng2023distillation ; yin2024one , these approaches often overlook the core architectural inefficiencies of DiT itself. Unlike U-Net-based architectures ronneberger2015u , which mitigate complexity through a contracting-expansive structure, DiT’s reliance on token-level self-attention incurs high costs that scale with model size and token count. This bottleneck highlights a critical need for DiT-specific innovations that manage token density intelligently, balancing efficiency and quality without sacrificing generation fidelity.
@@
To address this challenge, we examine the distribution of attention across DiT’s layers and sampling steps, seeking to uncover structural patterns that could guide more efficient token management. As illustrated in Figure 1(a), attention maps reveal that tokens capture different levels of granularity across layers: in the bottom layers, such as layer 0, attention maps approximately appear uniform distribution, indicating a focus on broad, global features akin to global pooling. Meanwhile, middle layers alternate in their attention, with certain layers (e.g., layer 6 and layer 26) capturing local details, while others (e.g., layer 22) emphasize global structure. This observation is quantified in Figure 1(b), where the variance in attention scores highlights DiT’s consistent pattern of alternation between global and local feature extraction across layers. In addition to spatial analysis, Figure 1(b) also reveals insights across sampling steps. We observe that, while the pattern of alternation between global and local focus remains stable, attention variance increases as the denoising process advances, indicating a growing emphasis on local information at lower-noise stages. This analysis yields three core insights: (1) the initial self-attention layers exhibit minimal variance, showing low discrepancy in feature extraction across tokens; (2) DiT’s architecture inherently alternates between global and local feature focus across layers, a pattern consistent across all sampling steps; and (3) as denoising progresses, the model increasingly prioritizes local details, dynamically adapting to heightened demands for detail at later stages.
@@
Building on these insights, we propose SparseDiT, which reimagines token density management as a dynamic process adapting across both spatial layers and temporal stages. Spatially, shallow layers capture smooth, global features, making complex self-attention less efficient. Subsequent layers alternate between high-frequency local details and low-frequency global information, where sparse tokens efficiently capture global features, and dense tokens refine local details. Temporally, as denoising advances, the need for local detail increases, motivating a dynamic token adjustment strategy. This dual-layered adaptation integrates layer-wise token modulation—balancing efficiency and detail within each sampling step—and a timestep-wise pruning strategy that adjusts token density dynamically. Together, these spatial and temporal adaptations allow SparseDiT to achieve computational efficiency without sacrificing high-fidelity detail, as reflected in SparseDiT’s architectural design and pruning strategy, tailored to meet the unique spatial and temporal demands of the generation process.
@@
SparseDiT Architecture: SparseDiT’s architecture employs a three-segment design that aligns token density with the features each layer captures. In the bottom layers, we replace self-attention with a poolingformer structure to capture broad global features through average pooling, reducing computation while preserving essential structure. In the middle layers, SparseDiT introduces Sparse-Dense Token Modules (SDTM), blending sparse tokens for global structure with dense tokens to refine local details. This design enables SparseDiT to retain efficiency while maintaining stability and detail. In the top layers, the model processes all tokens densely, focusing on high-frequency details for refined output quality.
@@
Timestep-Wise Pruning Rate Strategy: Complementing spatial adjustments, SparseDiT’s timestep-wise pruning rate strategy adapts token density across denoising stages. In early stages, where broad, low-frequency structures dominate, SparseDiT applies a high pruning rate, conserving resources. As the process progresses and intricate details emerge, the pruning rate decreases, allowing for a gradual increase in token density. This adjustment aligns computational effort with each stage’s complexity, dynamically balancing efficiency and detail to optimize resources for high-fidelity output.
@@
Our empirical results substantiate the effectiveness of SparseDiT’s integrated approach. SparseDiT reduces the FLOPs of DiT-XL by 55% and improves inference speed by 175% on $\times512$ ImageNet deng2009imagenet images, with only a increase of 0.09 in FID score heusel2017gans . On the Latte-XL dataset ma2024latte , SparseDiT achieves a 56% reduction in FLOPs across standard video generation datasets, including FaceForensics rossler2018faceforensics , SkyTimelapse xiong2018learning , UCF101 soomro2012dataset , and Taichi-HD siarohin2019first . Additionally, on the more challenging text-to-image generation task, we achieve a 69% improvement in inference speed on PixArt-$\alpha$ with a 0.24 FID score reduction. These results demonstrate SparseDiT’s capability to provide an efficient, high-quality architecture for diffusion-based generation, compatible with further sampling optimization techniques for enhanced efficiency.
(a) Architecture of SparseDiT.
@@
Figure 2: Architecture of SparseDiT and timestep-wise pruning rate strategy. The architecture consists of three segments: bottom, middle, and top. The bottom segment includes poolingformers (PF). The middle segment comprises multiple sparse-dense token modules (SDTMs), where sparse token generation transformers (“Sparse Gen") and dense token recovery transformers (“Dense Rec") alternate to balance global and detailed information processing. The top segment contains standard transformers (“Tr") for final processing. During the denoising process, we apply varying pruning rates $r$ at different stages (depicted as four stages).
@@ 2 Related works
This direction focuses on optimizing sampling steps in diffusion models. Some approaches song2020denoising ; lu2022dpm ; bao2022analyticdpm ; lu2022dpm++ design new solvers to achieve faster sampling with fewer steps. Consistency models (CMs) song2023consistency ; song2023improved ; luo2023latent ; wang2024phased ; kim2024consistency are closely related to diffusion models, achieving by distilling pre-trained diffusion models or training from scratch. These models learn a one-step mapping between noise and data, and all the points of the sampling trajectory map to the same initial data point. Flow-based lipman2022flow ; liu2022flow ; zhu2024flowie ; esser2024scaling ; zhu2025slimflow approaches straighten the transport trajectories among different distributions and thereby realize efficient inference. Knowledge distillation are applied by some methods song2023consistency ; salimans2022progressive ; luo2023latent ; meng2023distillation ; yin2024one to reduce the sampling steps and match the output of the combined conditional and unconditional models. Parallel sampling techniques shih2024parallel ; zheng2023fast ; tang2024accelerating employ Picard-Lindelöf iteration or Fourier neural operators for solving SDE/ODE.
@@
Another avenue focuses on improving inference time within a single model evaluation. Some methods attempt to compress the size of diffusion models via pruning fang2023structural or quantization li2023q ; shang2023post . Spectral Diffusion yang2023diffusion boosts structure design by incorporating frequency dynamics and priors. OMS-DPM liu2023oms proposes a model schedule that selects small models or large models at specific sampling steps to balance generation quality and inference speed. Deepcache ma2024deepcache , TokenCache lou2024token , and DiTFastAttnyuan2025ditfastattn reuse or share similar features across sampling steps. The early stopping mechanism in diffusion is explored in  li2023autodiffusion ; moon2023early ; tang2023deediff .
Most of the above methods are designed for general diffusion models. Recently, the Diffusion Transformer (DiT) has emerged as a more potential backbone, surpassing the previously dominant U-Net architectures. Unlike U-Net-based models, DiT’s inefficiency stems largely from the self-attention mechanism. Limited research has specifically addressed the efficiency of DiT-based models. Pu et al. pu2024efficient identifies the redundancies in the self-attention and adopts timestep dynamic mediator tokens to compress attention complexity, achieving some performances but limited FLOPs reduction, as this approach primarily reduces the number of keys and values rather than the overall tokens count. In Vision transformers for classification task, numerous methods wang2021not ; chen2021chasing ; fayyaz2022adaptive ; chang2023making ; bolya2023tome ; zong2022self ; meng2022adavit successfully removed redundant tokens to improve performance-efficiency trade-offs. ToMeSD bolya2023tomesd is the first to explore token reduction in diffusion models. However, its performance on DiT was suboptimal, as shown in  moon2023early ; zhao2024dynamic . U-Net can be regarded as a sparse network due to its contracting and expansive structure. EDT chen2024edt applies a novel architecture similar to U-Net, while they still have 1.4 FID score gap compared to its baseline model MDTv2 gao2023mdtv2 via 2, [^0] K iterations. These findings indicate that directly transferring token reduction techniques from classification tasks or reusing U-Net structures may be insufficient for DiT. Instead, a specialized token reduction strategy is needed. DyDiT zhao2024dynamic and DiffCR you2025layer design dynamic networks to compress DiT from multiple dimensions, such as token, layer, attention head, channel.
@@ 3 Method
We first introduce SparseDiT architecture in Section 3.1, followed by our timestep-wise pruning rate strategy in Section 5. Finally, Section 3.3 details the initialization and fine-tuning of our network.
3.1 SparseDiT
@@
The SparseDiT architecture is illustrated in Figure 2(a), where the transformer layers of DiT based model are divided into three groups: bottom, middle, and top. The bottom layers leverage poolingformers to efficiently capture global features. The middle layers contain multiple sparse-dense token modules (SDTM), which decouple the representation process into global structure capture and local details enhancement using sparse and dense tokens, respectively. The top layers retain the standard transformers, processing dense tokens to generate the final predictions at each sampling step. The primary computational savings are achieved through sparse tokens, hence the majority of transformer layers are located in the middle section to maximize efficiency.
@@
(b) Modifying first two attention in DiT-XL.
Figure 3: Comparison of images generated by the standard DiT-XL model and a modified DiT-XL model where the first two attention maps are replaced with full-one matrices without fine-tuning. Both sets of images exhibit similar visual quality.
@@
As shown in the first subfigure in Figure 1(a), attention scores in the bottom layers exhibit a nearly uniform distribution, with each token evenly extracting global features, akin to a global average pooling operator. To investigate further, we conducted a toy experiment where the attention maps in the first two transformer layers were replaced with matrices filled with a constant value (e.g., 1), without any fine-tuning. Figure 3 shows that, using the same initial noise and random seed, the images generated by the original DiT-XL and the modified DiT-XL are nearly identical, suggesting that complex self-attention calculations offer limited additional information. Given that self-attention in the bottom layers captures global features, we further question whether sparse tokens can be used in these layers. However, experiments (in Table 15) demonstrated that applying sparse tokens in the bottom layers results in unstable training, highlighting the necessity of retaining complete tokens in these layers.
Based on the above analysis, we adopt poolingformers to replace the original transformers. In poolingformer, we remove queries and keys as attention maps are not computed. Instead, we perform a global average pooling on the value $V\in\mathbb{R}^{N\times C}$ and integrate it into input tokens $X$, which can be represented as
@@
where $\bar{V}$ is the mean along the dimension of $N$, i.e., $\bar{v}=\frac{1}{N}\sum V(\bar{v}\in\mathbb{R}^{1\times C})$, and then repeating $N$ times to shape $N\times C$. The parameters from adaLN block are omitted in all the equations for brevity. Our poolingformer can be viewed as a special case of the model in yu2022metaformer , behaving identically when the pooling kernel size is equal to the input size.
@@
We present the sparse-dense token module (SDTM) to generate sparse and dense tokens processed by the corresponding transformer layers. The high-level idea is to decouple global structure extraction and local detail extraction. Sparse tokens capture global structure information and reduce computational cost, while dense tokens enhance local detail and stabilize training. Sparse and dense tokens are converted to each other within SDTM.
Initially, SDTM introduces a set of sparse tokens $X_{s}\in\mathbb{R}^{M\times C}$, where $M$ is the number of sparse tokens, typically $M\ll N$. These sparse tokens can be initialized by adaptively pooling the dense tokens $X$. We define pruning rate r=1M/Nr=1-M/N to represent the sparsity degree. To store the structure information, we first reshape the dense tokens $X\in\mathbb{R}^{N\times C}$ into the latent shape. For instance, if the input is an image, we reshape it into shape $H\times W\times C$, then pool it across the spatial dimensions to shape $H^{\prime}\times W^{\prime}\times C$, where $H^{\prime}\times W^{\prime}=M$. The intentions of spatial pooling initialization are two-fold. First, the initial sparse tokens can be distributed uniformly in space and the representation of each sparse token is associated with a specific spatial location, which is beneficial for downstream tasks such as image editing huang2024diffusion . Second, it can prevent the semantic tokens from collapsing to one point in the following layers. Sparse token interact with full-size dense tokens via an attention layer to integrate the global information:
@@
where the triplet input of $MHA$ are queries, keys, and values in turn.
The generated sparse tokens are fed into subsequent transformer layers, referred to as sparse transformers. Since $M\ll N$, our SDTM can substantially reduce computational cost.
Following this, we restore dense tokens from sparse tokens. First, we reshape the sparse tokens $X_{s}\in\mathbb{R}^{M\times C}$ into structure shape $H^{\prime}\times W^{\prime}\times C$ and upsample it to the same shape as input dense tokens $X\in\mathbb{R}^{H\times W\times C}$. We introduce two linear layers to combine upsampling sparse tokens with dense tokens, which is represented as:
@@
At the end of SDTM, several transformer layers, termed dense transformers, process full-size dense tokens to enhance local details.
@@
Table 1: Comparison of performance between SparseDiT and DiT for class-conditional image generation task on ImageNet.
3.2 Timestep-wise pruning rate